OpenStudy (caozeyuan):

Suppose A is aa 2x2 matrix with real number entries, FInd such A that A^2=-I, where I is the 2x2 identity matrix @amistre64

OpenStudy (caozeyuan):

@RosieS12 Hey there, any idea for this one?

OpenStudy (anonymous):

well i think we have to find what x and y equal and then plug them in the equation?

OpenStudy (caozeyuan):

what do you mean x and y?

OpenStudy (anonymous):

oh im really sorry i was looking at the wrong question xc my bad :)

OpenStudy (amistre64):

a b a b -1 0 c d a b = 0 -1 aa + ab = -1 ac + ad = 0 ba + db = 0 bc + dd = -1

OpenStudy (amistre64):

you essentially create 2 sets of systems of equations to solves

OpenStudy (amistre64):

opps, i typoed myself

OpenStudy (caozeyuan):

I didn't understand your first response, although I trust you one the accuracy

OpenStudy (amistre64):

a b a b -1 0 c d c d = 0 -1 aa + cb = -1 ac + cd = 0 ba + db = 0 bc + dd = -1

OpenStudy (amistre64):

or we can augment A with -I, and then equate the results back to A

OpenStudy (caozeyuan):

but these equations contain 4 variables and they are all second-order, isn't too complicated?

OpenStudy (amistre64):

its not too complicated

OpenStudy (caozeyuan):

why augment A with -I, how does that help

OpenStudy (caozeyuan):

but it's too complicated for both my self and my TI-NSpire, maybe it isn't for you genius. LOL

OpenStudy (caozeyuan):

@Loser66 , I don't know the question just said that

OpenStudy (caozeyuan):

NO here is the answer form the back of my book